# On the discontinuities of Hausdorff dimension in generic dynamical Lagrange spectrum

### v1

@@ 3.2. A technical proposition
We will now give an explicit positive lower bound for $\delta$ in terms of $\eta$. In order to do that, we define recursively, for each integer $n\geq 0$ and $x\in\mathbb{R}$, the function $\mathcal{T}(n,x)$ by $\mathcal{T}(x,0)=x$, $\mathcal{T}(x,n+1)=e^{\mathcal{T}(x,n)}$. We have for $n\geq 0$
@@
since $p_{n}\geq p_{0}>\lceil 2/\tau^{2}\rceil$. Therefore $r_{0}=p_{s_{0}}<\mathcal{T}(p_{0},2s_{0})$ and
@@
As $p_{0}=\max\{\lceil\frac{40000(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\eta^{2}}\rceil,z(\eta,\Lambda)\}$ and $s_{0}<(1+\frac{2}{\tau})\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}=(1+\frac{200}{\eta})\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}$, we have by 3.7
$\delta>\frac{c_{3}e^{-c_{1}}}{\mathcal{T}(p_{0},2s_{0}+3)}=\frac{c_{3}e^{-c_{1}}}{\mathcal{T}(\max\{\lceil\frac{40000(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\eta^{2}}\rceil,z(\eta,\Lambda)\},\lceil\frac{201}{\eta}\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}\rceil)},$
@@
Given $\epsilon>0$ and $c_{0}>0$ there exists a constant $\delta=\delta(\epsilon,c_{0})>0$ such that if $X$ is a compact $\varphi$-invariant subset of $\Lambda$ such that the limit capacities $D_{u}(X)$ and $D_{s}(X)$ satisfy both $D_{u}(X),D_{s}(X)\geq c_{0}$. Then there are subhorseshoes $\Lambda^{s}(X)$ and $\Lambda^{u}(X)$ of $\Lambda$ such that
@@ 3.3. First accumulation point of the Lagrange spectrum
Let $c^{*}_{\varphi,f}=\min L^{{}^{\prime}}_{\varphi,f}$. Given $\epsilon>0$, it is clearly that $L_{\varphi,f}(c^{*}_{\varphi,f}-\epsilon)=0$ because $\mathcal{L}_{\varphi,f}\cap(-\infty,c^{*}_{\varphi,f}-\epsilon)$ is finite. On the other hand, take an injective sequence $(y_{n})_{n\in\mathbb{N}}=(\ell_{\varphi,f}(x_{n}))_{n\in\mathbb{N}}\subset\mathcal{L}_{\varphi,f}$ such that $\lim\limits_{n\rightarrow\infty}y_{n}=c^{*}_{\varphi,f}$ and consider $N\in\mathbb{N}$ big enough such that for two elements $x,y\in\Lambda$ if their kneading sequences coincide in the central block (centered at the zero position) of size $12N+$ 1 then $\lvert f(x)-f(y)\rvert<\epsilon/6$.
Take first $n_{0}\in\mathbb{N}$ large so that $\lvert\ell_{\varphi,f}(x_{n})-c^{*}_{\varphi,f}\rvert<\epsilon/6$ for $n\geq n_{0}$ and there are infinitely many $j\in\mathbb{N}$ such that $\lvert f(\varphi^{j}(x_{n}))-c^{*}_{\varphi,f}\rvert<\epsilon/6$. Given such a pair $(j,n)$, consider the finite sequence with $12N+$ 1 terms $S(j,n)=(b_{j-N}^{(n)},b_{j-N+1}^{(n)},\cdots,b_{j}^{(n)},\cdots,b_{j+N}^{(n)})$ where $\Pi^{-1}((b_{j}^{(n)})_{j\in\mathbb{Z}})=x_{n}$. There is a sequence $S$ such that for infinitely many values of $n$, $S$ appears infinitely many times as $S(j,n)$; i.e., there are $j_{1}(n)<j_{2}(n)<\cdots$ with $\lim\limits_{i\rightarrow\infty}(j_{i+1}(n)-j_{i}(n))=\infty$ and $S(j_{i}(n),n)=S$ for all $i\geq 1$ and for all $n$ in some infinite set $A\subset\mathbb{N}$.
@@
we have $\Sigma(C)$ is a complete subshift and for $x\in\Lambda(\Sigma(C))=\Lambda_{C}$ (the subhorseshoe associated to $\Sigma(C)$) we have $m_{\varphi,f}(x)<c^{*}_{\varphi,f}+\epsilon/2$. Indeed, for every $k\in\mathbb{Z}$ the kneading sequence of $\varphi^{k}(x)$ coincides in the central block of size $12N+$ 1 with the kneading sequence of $\varphi^{l}(x_{\theta})$ where $\theta$ is either $n_{1}$ or $n_{2}$ and $l\geq r$. So
@@
It follows from the previous proposition and 3.8 that $0<L_{\varphi,f}(c_{\varphi,f}+\epsilon)\leq HD(\Lambda_{c_{\varphi,f}+\epsilon}).$ Now, if $HD(\Lambda_{c_{\varphi,f}})>0$ then by 2.7, $D_{u}(\Lambda_{c_{\varphi,f}})>0$ (also $D_{s}(\Lambda_{c_{\varphi,f}})>0$), and by proposition 3.4 we can find some horseshoe $\tilde{\Lambda}\subset\Lambda_{c_{\varphi,f}-\delta}$ for some $\delta>0$ and arguing as before, we get the contradiction $L_{\varphi,f}(c_{\varphi,f}-\delta/2)>0.$
@@
This corollary remains true if $HD(\Lambda)\geq 1$ because proposition 1 of [9] let us also show the existence of $\tilde{\Lambda}$ and $\delta>0$ as before.
@@
Suppose $\lim\limits_{t\to\ c_{\varphi,f}^{+}}HD(\Lambda_{t})=h>0$, then by 2.7, for $t>c_{\varphi,f}$ one has $D_{u}(\Lambda_{t})\geq h/(1+\tilde{C}).$ On the other hand, proposition 3.4 let us find some $\delta=\delta(\frac{1}{2},\frac{h}{1+\tilde{C}})>0$ such that for any $t>c_{\varphi,f}$ we can find some horseshoe $\Lambda^{u}(\Lambda_{t})\subset\Lambda_{t-\delta}$ (the other conclusions of the proposition are not necessary here). By applying this to $t=c_{\varphi,f}+\delta/2$, we get the contradiction $0<HD(\Lambda^{u}(\Lambda_{c_{\varphi,f}+\delta/2}))\leq HD(\Lambda_{c_{\varphi,f}-\delta/2})$. Then
@@
This corollary also holds when $HD(\Lambda)\geq 1$ because as we will see later, before $\tilde{c}_{\varphi,f}$, it is true some expression of the type $L_{\varphi,f}=\max\limits_{i}L_{i}$, where the functions $L_{i}$ are defined like $L_{\varphi,f}$ but are associated to horseshoes with Hausdorff dimension less than $1$.
@@ 3.4. Geometric consequences of having a discontinuity
Take $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and suppose $t_{0}\in\mathbb{R}$ is a discontinuity of the map $t\rightarrow L_{\varphi,f}(t)=HD(\mathcal{L}_{\varphi,f}\cap(-\infty,t))$. So, there exists an $a>0$ such that
@@
By corollary 3.10 and 3.8 we have $0<L_{\varphi,f}(t_{0})\leq HD(\Lambda_{t_{0}}),$ then $D_{u}(\Lambda_{t_{0}})>0$ and one more time, by proposition 3.4, we can find some horseshoe $\Lambda^{0}\subset\Lambda_{t_{0}}$. For $20<\epsilon<a/$ 2 and $c_{0}=HD(\Lambda^{0})/(\tilde{C}+1)>0$ take $\delta=\delta(\epsilon/2k,c_{0})<\epsilon$ as in the corollary 3.5 where $k>1$ is a Lipschitz’s constant for $f$, and let us consider for $t\in\mathbb{R}$ and $n\in\mathbb{N}$ the set $C(\Lambda_{t},n)$. By compactness, one has
@@
Take then, $n=n(\delta/2k)$ and consider the maximal invariant set
@@
for $t_{0}<t\leq t(n)$.
Observe that for $x\in P$ and $m\in\mathbb{Z}$ if $y\in\Lambda_{t_{0}}$ belongs to the same rectangle $R(\theta;0)$ as $\varphi^{m}(x)$ for some $\theta\in C(\Lambda_{t_{0}},n)$ then
@@
Now, by proposition 2.3, the set $P$ admits a decomposition $P=\bigcup\limits_{i\in\mathcal{I}}\tilde{\Lambda}_{i}$ where $\mathcal{I}$ is a finite index set and for any $i\in\mathcal{I}$$\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set. In particular, given $i_{1}\in\mathcal{I}$ we can find $i_{2}\in\mathcal{I}$ such that $\tilde{\Lambda}_{i_{2}}$ is a subhorseshoe with $\omega(x)\subset\tilde{\Lambda}_{i_{2}}$ for every $x\in\tilde{\Lambda}_{i_{1}}$; and from this and 3.10, it follows that $\ell_{\varphi,f}(x)=\ell_{\varphi,f}(y)$ for some $y\in\tilde{\Lambda}_{i_{2}}$. We conclude then
@@
Let $\tilde{\Lambda}_{i_{0}}$ with $HD(\ell_{\varphi,f}(P))=HD(\tilde{\Lambda}_{i_{0}})$. As $\Lambda^{0}\subset P$, by 2.6 and 2.7 one has
@@
Now, suppose there exists a subhorseshoe $\widetilde{\Lambda}\subset\Lambda_{q}$ for some $q<t_{0}$ with $\Lambda^{u}(t_{0})\cup\Lambda^{s}(t_{0})\subset\widetilde{\Lambda}$, then as $\Lambda_{t}\subset P$ for $t_{0}<t\leq t(h)$, we have by 3.9 and lemma 3.6
@@
which is a contradiction. Then, by definition, $\Lambda^{s}(t_{0})$ and $\Lambda^{u}(t_{0})$ don’t connect before $t_{0}$.
On the other hand, fix $x\in\Lambda^{s}(t_{0})$, $y\in\Lambda^{u}(t_{0})$ with kneading sequences $(x_{n})_{n\in\mathbb{Z}}$, respectively $(y_{n})_{n\in\mathbb{Z}}$. As the sets $(\tilde{\Lambda}_{i_{0}})_{\epsilon/2k}(x)$ and $(\tilde{\Lambda}_{i_{0}})_{\epsilon/2k}(y)$ are nonempty we can find two words $\theta$ and $\tilde{\theta}$ in $C(\tilde{\Lambda}_{i_{0}},n(\epsilon/2k))$ that appear respectively in the sequences $(x_{n})_{n\in\mathbb{Z}}$ and $(y_{n})_{n\in\mathbb{Z}}$ as sub-words and also appear in the kneading sequence of two points $\tilde{x}_{1},\tilde{y}_{1}\in\tilde{\Lambda}_{i_{0}}$, i.e., $\tilde{x}_{1}\in R(\theta;0)$, and $\tilde{y}_{1}\in R(\tilde{\theta};0)$, $(x_{N_{1}},\dots x_{N_{1}-\lvert\theta\rvert-1})=\theta$ and $(y_{-N_{2}-\lvert\tilde{\theta}\rvert+1},\dots y_{-N_{2}})=\tilde{\theta}$ for some $N_{1},N_{2}>0$.
@@
Analogously we can find $\tilde{z}\in W^{u}(\Lambda^{u}(t_{0}))\cap W^{s}(\Lambda^{s}(t_{0}))\cap\tilde{P}$. Moreover, as $\Lambda^{u}(t_{0})\cup\Lambda^{s}(t_{0})\cup\tilde{\Lambda}_{i_{0}}\subset\Lambda_{t_{0}+\delta/2}$, reasoning as we did for $P$, we have $\tilde{P}\subset\Lambda_{k\cdot\epsilon/2k+t_{0}+\delta/2}=\Lambda_{\epsilon/2+t_{0}+\delta/2}.$ That is,
@@
Take $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$$f\in\mathcal{P}_{\varphi,\Lambda}$ and some discontinuity $t_{0}$ of the map
@@
  • $\Lambda^{s}(t_{0})$ and $\Lambda^{u}(t_{0})$ don’t connect before $t_{0}$,
@@
As in remark 3.11, this result also holds when $HD(\Lambda)\geq 1$ and $t_{0}<c_{\varphi,f}$. Note that, in our context, by corollary 3.3, $L_{\varphi,f}$ is discontinuous in $\tilde{c}_{\varphi,f}$ if and only if $L_{\varphi,f}(\tilde{c}_{\varphi,f})<1.$

### v2

@@ 3.2. A technical proposition
We will now give an explicit positive lower bound for $\delta$ in terms of $\eta$. In order to do that, we define recursively, for each integer $n\geq 0$ and $x\in\mathbb{R}$, the function $\mathcal{T}(n,x)$ by $\mathcal{T}(x,0)=x$, $\mathcal{T}(x,n+1)=e^{\mathcal{T}(x,n)}$. By (2.3), we have for $n\geq 0$
@@
since $p_{n}\geq p_{0}>\lceil 2/\tau^{2}\rceil$ and $p_{n}>1/\kappa^{2}$. Therefore $r_{0}=p_{s_{0}}<\mathcal{T}(p_{0},2s_{0})$ and
@@
As $p_{0}=\max\{\lceil\frac{40000(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\eta^{2}}\rceil,r(\eta,\Lambda)\}$ and $s_{0}<(1+\frac{2}{\tau})\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}=(1+\frac{200}{\eta})\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}$, we have by (3.9)
$\delta>\frac{c_{3}e^{-c_{1}}}{\mathcal{T}(p_{0},2s_{0}+3)}=\frac{c_{3}e^{-c_{1}}}{\mathcal{T}(\max\{\lceil\frac{40000(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\eta^{2}}\rceil,r(\eta,\Lambda)\},\lceil\frac{201}{\eta}\log\frac{4(\alpha_{1}+\alpha_{2}+1)}{\eta}\rceil)},$
@@
Given $\epsilon>0$ and $c_{0}>0$ there exists a constant $\delta=\delta(\epsilon,c_{0})>0$ such that if $X$ is a compact $\varphi$-invariant subset of $\Lambda$ such that the limit capacities $D_{u}(X)$ and $D_{s}(X)$ satisfy both $D_{u}(X),D_{s}(X)\geq c_{0}$, then there are subhorseshoes $\Lambda^{s}(X)$ and $\Lambda^{u}(X)$ of $\Lambda$ such that
@@ 3.3. First accumulation point of the Lagrange spectrum
Let $c^{*}_{\varphi,f}=\min L^{{}^{\prime}}_{\varphi,f}$. Given $\epsilon>0$, it is clear that $L_{\varphi,f}(c^{*}_{\varphi,f}-\epsilon)=0$ because $\mathcal{L}_{\varphi,f}\cap(-\infty,c^{*}_{\varphi,f}-\epsilon)$ is finite. On the other hand, take an injective sequence $(y_{n})_{n\in\mathbb{N}}=(\ell_{\varphi,f}(x_{n}))_{n\in\mathbb{N}}\subset\mathcal{L}_{\varphi,f}$ such that $\lim\limits_{n\rightarrow\infty}y_{n}=c^{*}_{\varphi,f}$ and consider $N\in\mathbb{N}$ big enough such that for two elements $x,y\in\Lambda$ if their kneading sequences coincide in the central block (centered at the zero position) of size $2N+1$ then $\lvert f(x)-f(y)\rvert<\epsilon/6$.
Take first $n_{0}\in\mathbb{N}$ large so that $\lvert\ell_{\varphi,f}(x_{n})-c^{*}_{\varphi,f}\rvert<\epsilon/6$ for $n\geq n_{0}$ and there are infinitely many $j\in\mathbb{N}$ such that $\lvert f(\varphi^{j}(x_{n}))-c^{*}_{\varphi,f}\rvert<\epsilon/6$. Given such a pair $(j,n)$, consider the finite sequence with $2N+1$ terms $S(j,n)=(b_{j-N}^{(n)},b_{j-N+1}^{(n)},\cdots,b_{j}^{(n)},\cdots,b_{j+N}^{(n)})$ where $\Pi^{-1}((b_{j}^{(n)})_{j\in\mathbb{Z}})=x_{n}$. There is a sequence $S$ such that for infinitely many values of $n$, $S$ appears infinitely many times as $S(j,n)$; i.e., there are $j_{1}(n)<j_{2}(n)<\cdots$ with $\lim\limits_{i\rightarrow\infty}(j_{i+1}(n)-j_{i}(n))=\infty$ and $S(j_{i}(n),n)=S$ for all $i\geq 1$ and for all $n$ in some infinite set $A\subset\mathbb{N}$.
@@
we have $\Sigma(C)$ is a complete subshift and for $x\in\Lambda(\Sigma(C))=\Lambda_{C}$ (the subhorseshoe associated to $\Sigma(C)$) we have $m_{\varphi,f}(x)<c^{*}_{\varphi,f}+\epsilon/2$. Indeed, for every $k\in\mathbb{Z}$, the kneading sequence of $\varphi^{k}(x)$ coincides in the central block of size $2N+1$ with the kneading sequence of $\varphi^{l}(x_{\theta})$ where $\theta$ is either $n_{1}$ or $n_{2}$ and $l\geq r$. So,
@@
It follows from the previous proposition and (3.10) that $0<L_{\varphi,f}(c_{\varphi,f}+\epsilon)\leq HD(\Lambda_{c_{\varphi,f}+\epsilon}).$ Now, if $HD(\Lambda_{c_{\varphi,f}})>0$ then by (2.10), $D_{u}(\Lambda_{c_{\varphi,f}})>0$ (also $D_{s}(\Lambda_{c_{\varphi,f}})>0$), and by Proposition 3.4 we can find some horseshoe $\tilde{\Lambda}\subset\Lambda_{c_{\varphi,f}-\delta}$ for some $\delta>0$ and arguing as before, we get the contradiction $L_{\varphi,f}(c_{\varphi,f}-\delta/2)>0.$
@@
This corollary remains true if $HD(\Lambda)\geq 1$ because of [9, Proposition 1] lets us also show the existence of $\tilde{\Lambda}$ and $\delta>0$ as before.
@@
[…]
@@
This corollary also holds when $HD(\Lambda)\geq 1$ because as we will see later, before $\tilde{c}_{\varphi,f}$, it is true some expression of the type $L_{\varphi,f}=\max\limits_{i}L_{i}$, where the functions $L_{i}$ are defined as $L_{\varphi,f}$ but are associated to horseshoes with Hausdorff dimension less than $1$.
@@ 3.4. Geometric consequences of having a discontinuity
Take $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and suppose $t_{0}\in\mathbb{R}$ is a discontinuity of the map $t\mapsto L_{\varphi,f}(t)=HD(\mathcal{L}_{\varphi,f}\cap(-\infty,t))$. So, there exists an $a>0$ such that
@@
By Corollary 3.10 and (3.10) we have $0<L_{\varphi,f}(t_{0})\leq HD(\Lambda_{t_{0}}),$ then $D_{u}(\Lambda_{t_{0}})>0$ and by Proposition 3.4, we can find some horseshoe $\Lambda^{0}\subset\Lambda_{t_{0}}$. For $0<\epsilon<a/2$ and $c_{0}=HD(\Lambda^{0})/(\tilde{C}+1)>0$, where $\tilde{C}$ si given in (2.7), take $\delta=\delta(\epsilon/2k,c_{0})<\epsilon$ as in Corollary 3.5 where $k>1$ is a Lipschitz’s constant for $f$.
Let us consider for $t\in\mathbb{R}$ and $n\in\mathbb{N}$ the set $C(\Lambda_{t},n)$, defined in Subsection 3.2. By compactness, one has
@@
Take then, $n=n(\delta/2k)=\min\{r\in\mathbb{N}:\forall\theta\in C(\Lambda,r),\ \text{diam}(R(\theta;0))\leq\delta/4k\}$ and consider the maximal invariant set
@@
for $t_{0}<t\leq t(n)$. Observe that given $x\in P$ and $m\in\mathbb{Z}$ if $y\in\Lambda_{t_{0}}$ belongs to the same rectangle $R(\theta;0)$ as $\varphi^{m}(x)$ for some $\theta\in C(\Lambda_{t_{0}},n)$ then
@@
Now, by Proposition 2.3, the set $P$ admits a decomposition $P=\bigcup\limits_{i\in\mathcal{I}}\tilde{\Lambda}_{i}$ where $\mathcal{I}$ is a finite index set and for any $i\in\mathcal{I}$, $\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set. In particular, given $i_{1}\in\mathcal{I}$ we can find $i_{2}\in\mathcal{I}$ such that $\tilde{\Lambda}_{i_{2}}$ is a subhorseshoe with $\omega(x)\subset\tilde{\Lambda}_{i_{2}}$ for every $x\in\tilde{\Lambda}_{i_{1}}$; and from this and (3.12), it follows that $\ell_{\varphi,f}(x)=\ell_{\varphi,f}(y)$ for some $y\in\tilde{\Lambda}_{i_{2}}$. We conclude then
@@
Let $\tilde{\Lambda}_{i_{0}}$ with $HD(\ell_{\varphi,f}(P))=HD(\tilde{\Lambda}_{i_{0}})$. As $\Lambda^{0}\subset P$, by (2.9) and (2.10) one has
@@
Now, suppose there exists a subhorseshoe $\widetilde{\Lambda}\subset\Lambda_{q}$ for some $q<t_{0}$ with $\Lambda^{u}(t_{0})\cup\Lambda^{s}(t_{0})\subset\widetilde{\Lambda}$, then as $\Lambda_{t}\subset P$ for $t_{0}<t\leq t(h)$, we have by (3.11) and Lemma 3.6
@@
which is a contradiction. Then, by definition, $\Lambda^{s}(t_{0})$ and $\Lambda^{u}(t_{0})$ do not connect before $t_{0}$.
On the other hand, fix $x\in\Lambda^{s}(t_{0})$, $y\in\Lambda^{u}(t_{0})$ with kneading sequences $(x_{n})_{n\in\mathbb{Z}}$, respectively $(y_{n})_{n\in\mathbb{Z}}$. As the sets $(\tilde{\Lambda}_{i_{0}})_{\epsilon/2k}(x)$ and $(\tilde{\Lambda}_{i_{0}})_{\epsilon/2k}(y)$ are nonempty, we can find two words $\theta$ and $\tilde{\theta}$ in $C(\tilde{\Lambda}_{i_{0}},n(\epsilon/2k))$ that appear respectively in the sequences $(x_{n})_{n\in\mathbb{Z}}$ and $(y_{n})_{n\in\mathbb{Z}}$ as sub-words and also appear in the kneading sequence of two points $\tilde{x}_{1},\tilde{y}_{1}\in\tilde{\Lambda}_{i_{0}}$, i.e., $\tilde{x}_{1}\in R(\theta;0)$, and $\tilde{y}_{1}\in R(\tilde{\theta};0)$, $(x_{N_{1}},\dots x_{N_{1}+\lvert\theta\rvert-1})=\theta$ and $(y_{-N_{2}-\lvert\tilde{\theta}\rvert+1},\dots y_{-N_{2}})=\tilde{\theta}$ for some $N_{1},N_{2}>0$.
@@
Analogously, we can find $\tilde{z}\in W^{u}(\Lambda^{u}(t_{0}))\cap W^{s}(\Lambda^{s}(t_{0}))\cap\tilde{P}$. Moreover, as $\Lambda^{u}(t_{0})\cup\Lambda^{s}(t_{0})\cup\tilde{\Lambda}_{i_{0}}\subset\Lambda_{t_{0}+\delta/2}$, reasoning as we did for $P$, we have $\tilde{P}\subset\Lambda_{k\cdot\epsilon/2k+t_{0}+\delta/2}=\Lambda_{\epsilon/2+t_{0}+\delta/2}.$ That is,
@@
Take $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and some discontinuity $t_{0}$ of the map
@@
  • $\Lambda^{s}(t_{0})$ and $\Lambda^{u}(t_{0})$ do not connect before $t_{0}$,
@@
As in Remark 3.11, this result also holds when $HD(\Lambda)\geq 1$ and $t_{0}<\tilde{c}_{\varphi,f}$. Note that, in our context, by Corollary 3.3, $L_{\varphi,f}$ is discontinuous in $\tilde{c}_{\varphi,f}$ if and only if $L_{\varphi,f}(\tilde{c}_{\varphi,f})<1.$